OpenStudy (anonymous):

What is the solution of the system? 2x - y = -7 4x - y = -4 Can someone show me how?

OpenStudy (anonymous):

are you trying to solve for x?

OpenStudy (freckles):

like notice the y in both equations have the same sign multiply one of the equations by -1 then add the equations together this way the y's will be eliminated and you can solve for x

OpenStudy (anonymous):

Oh okay I understand. I was trying to do something else. I feel dumb now. Thanks so much freckles!

OpenStudy (freckles):

np

OpenStudy (triciaal):

another approach 2x + 7 = y from eqn 1 and 4x +4 = y from eqn 2 set the y;s equal 2x + 7 = 4x + 4 solve for x -2x = -3 x = 3/2 substitute to get y value using either eqn y = 2(3/2) + 7 = 10 y = 4(3/2)+ 4 = 6 + 4 = 10

OpenStudy (anonymous):

Thanks @triciaal I am going to write that down.